# Syntriever: How to Train Your Retriever with Synthetic Data from LLMs

### v2

@@ Abstract
Syntriever: How to Train Your Retriever with Synthetic Data from LLMs
@@ 2.1 Problem Statement and Notation
where $\mathcal{M}$($\cdot$) denotes the LLM operation, and $\mathcal{P}_{\mathrm{cot}}$ denotes the prompt operator to generate CoT (see Appendix E for prompt details). $q^{\mathrm{cot}}_{i}$ contains sub-queries relevant to the original query, which are carefully planned out with clarification and details necessary to retrieve relevant documents. We will use $q^{\mathrm{cot}}_{i}$ as a positive document for $q_{i}$. This helps the retriever with understanding diverse contexts associated with related queries in the future.
@@
The derivation of (12) is provided in Appendix D. Thus, if we set $q=q_{i}$, $y_{\pi(1)}=c_{i}^{+}$, $y_{\pi(2)}=c_{i}^{-}$ and the rest of $y$’s as in-batch samples with $M=|B|$ and the reward $r(\cdot,\cdot)$ as the similarity metric $s_{\tau}(\cdot,\cdot)$, then (12) models the partial relation (10).
@@ 3.1 Experimental Settings
Settings of Syntriever. Syntriever uses pre-trained E5 Wang et al. (2022) as the base encoder $E$. In the settings of supervised fine-tuning, Syntriever is trained with synthetic data generated from each training dataset. In the settings of the zero-shot transfer, Syntriever is first trained on synthetic data based on training datasets of MSMARCO and NQ, and then is evaluated on out-of-domain datasets from BeIR benchmarks. This is a similar setting as Contriever Izacard et al. (2021), E5 Wang et al. (2022), etc. To minimize the effects of model sizes on performance, we set the size of Syntriever and all baseline models to (approximately) 125M. In alignment stage of Syntriever, we set $K=5$ by default, and set $N={K\choose 2}=10$. Detailed hyperparameters are in Appendix F.
@@
4 Experimental Results
Main Results. We first present the supervised fine-tuning results on seven datasets for the retrieval task which are BeIR benchmarks with training datasets. The results are shown in Table 1. Compared to the second-best models, Syntriever improves the retrieval performances by: 18.6% on MSMARCO, 5.9% on HotpotQA, 2.5% on FiQA, 1.8% on SciFact, 8.3% on NFCorpus, and 4% on NQ. The base encoder for Syntriever is a pre-trained E5; still, Syntriever achieves performance gain over E5 by: 18.6% on MSMARCO, 10.7% on HotpotQA, 10.6% on FiQA, 9.2% on SciFact, 20.9% on NFCorpus, 5.6% on Fever and 5.7% on NQ. This shows that Syntriever can successfully distill LLMs’ capability into small retrievers and improve their performance by a large margin. Overall, Syntriever shows robust performances on datasets both in generalized and specialized domains. Our results show that small LMs can efficiently learn from the teacher model through synthetic data and can be successfully aligned through feedback, even without access to the output probability of black-box LLMs.
@@ 4 Experimental Results
Ablation study. We conduct an ablation study on Syntriever. We add or remove model components, and the effects on the performance are shown in Table 3. The results show that both synthesized query $(q^{\text{cot}})$ and passages $(p^{+},p^{-})$ in the distillation stage improve the retrieval performances. Overall, the distillation stage achieves an average gain of 8.2% over the base retriever. Results show that the retriever successfully learns from the parametric knowledge of LLMs during the distillation stage. Also, the alignment component $(c^{+}\succ c^{-})$ in Table 3 is shown to achieve performance gains of up to 8.8%. Our results show that the alignment component is significant for retrieval performance, considering that nDCG@K is sensitive to the fine-grained ranking of relevant passages.
@@
Performances with different encoders. Syntriever is a framework for training encoders for retrieval, and thus can be combined with different sentence encoders. We experiment with various well-known encoders, e.g., ColBERT, SBERT, and Contriever, as the base encoders for Syntriever. Syntriever improves the performance by a large margin in all three retrieval models. The performance improvement is particularly high in nDCG@1 which concerns retrieving the exact passage relevant to the query. This is because the alignment stage in Syntriever helps the retriever with a fine-grained ranking of highly relevant passages. Overall, the results show that Syntriever is generally applicable to, and improves the performances of, various retrievers.
@@
Effects of Re-labeling Hallucination Passages. Table 5 shows that LLM self-verification and re-labeling are effective for the synthetic training by Syntriever. The performance improvement of self-verification on HotpotQA is relatively greater than other datasets. We found that, approximately 15% of synthetic positive passages were relabeled as hallucinations in the case of HotpotQA, whereas the proportion was about a few percent in other datasets. This indicates that the performance improvement through relabeling is likely higher for HotpotQA. In conclusion, removing hallucinations (and even re-using them as hard-negatives as in Syntriever) through self-verification is important for data synthesis, which is the case for most tasks utilizing LLMs Weng et al. (2023).
@@
Weaker but Cheaper LLMs can be effective. We examine how the LLM capabilities affect distillation and alignment performances. We consider two LLMs: GPT-4o vs. GPT-4o-mini, where GPT-4o is the larger and more capable model. First, we compare the distillation capabilities of two LLMs. Table 6 shows the comparison, where Syntriever is trained only up to the distillation stage. Interestingly, the distillation performance of GPT-4o-mini is better than GPT-4o on NFCorpus. Considering the datasets concern different knowledge domains (SciFact: scientific, NFCorpus: medical), smaller models may be better at teaching than larger ones in certain domains. Our results interestingly coincide with recent findings that weaker models may be better at teaching than stronger models in domains like math problem solving Bansal et al. (2024). Next, we examine the alignment capabilities of LLMs. For a fair comparison, two models are first trained by GPT-4o in the distillation stage, and then trained by different LLMs in the alignment stage. Table 7 shows that the larger model (GPT-4o) is better at alignment. It is challenging to rank top-$K$ passages retrieved by a distilled retriever, requiring a deep understanding of various contexts, and thus larger models may be more favored for the task. Overall, smaller models appear to be quite competitive, i.e., the performance gap is small or even better in some domains. Thus, our prospect is that distillation/alignment through small models will become an increasingly good alternative, especially under a fixed compute budget Bansal et al. (2024).
@@ Appendix A Reproducibility Statement
Black-box LLMs. We experiment with GPT-4o and GPT-4o-mini for synthetic data generation. Those models are accessible by OpenAI API. We generate LLMs’ responses using our prompt templates in Appendix E.
@@
Appendix B Additional Experiments
Method FiQA NFCorpus SciFact
Partial Plackett Luce 41.9 43.3 80.5
Bradley Terry 35.7 36.1 79.8
Table 8: Comparison of preference modeling. Results are in nDCG@10.
B.1 Comparison of Preference Modeling Methods
B.2 Effects of the number of retrieved passages during alignment
Appendix C Quality of Synthetic Positives
Passages Can derive answer Cannot derive answer
Synthetic positive 88% 12%
Ground truth 84% 16%
Appendix D Derivation of (12)
@@
Appendix E Prompt templates
E.1 Prompt template of positive passage generation $(\mathcal{P}_{+})$.
@@ Appendix E Prompt templates
E.2 Prompts for generating plausible but irrelevant passages $(\mathcal{P}_{-})$.
@@
E.3 Prompt template of relabeling for synthetic passages $(\mathcal{P}_{\text{Relabel}})$.
@@
E.4 Prompts for pair-wise comparison of two passages $(\mathcal{P}_{\text{Compare}})$.
@@ Appendix F Hyperparameters
Appendix G Dataset Statistics

### v3

# Syntriever: How to Train Your Retriever [Uncaptioned image] with Synthetic Data from LLMs

@@ Abstract
Syntriever: How to Train Your Retriever [Uncaptioned image] with Synthetic Data from LLMs
@@ 2.1 Problem Statement and Notation
where $\mathcal{M}$($\cdot$) denotes the LLM operation, and $\mathcal{P}_{\mathrm{cot}}$ denotes the prompt operator to generate CoT (see Appendix C for prompt details). $q^{\mathrm{cot}}_{i}$ contains sub-queries relevant to the original query, which are carefully planned out with clarification and details necessary to retrieve relevant documents. We will use $q^{\mathrm{cot}}_{i}$ as a positive document for $q_{i}$. This helps the retriever with understanding diverse contexts associated with related queries in the future.
@@
The derivation of (12) is provided in Appendix B. Thus, if we set $q=q_{i}$, $y_{\pi(1)}=c_{i}^{+}$, $y_{\pi(2)}=c_{i}^{-}$ and the rest of $y$’s as in-batch samples with $M=|B|$ and the reward $r(\cdot,\cdot)$ as the similarity metric $s_{\tau}(\cdot,\cdot)$, then (12) models the partial relation (10).
@@ 3.1 Experimental Settings
Settings of Syntriever. Syntriever uses pre-trained E5 Wang et al. (2022) as the base encoder $E$. In the settings of supervised fine-tuning, Syntriever is trained with synthetic data generated from each training dataset. In the settings of the zero-shot transfer, Syntriever is first trained on synthetic data based on training datasets of MSMARCO and NQ, and then is evaluated on out-of-domain datasets from BeIR benchmarks. This is a similar setting as Contriever Izacard et al. (2021), E5 Wang et al. (2022), etc. To minimize the effects of model sizes on performance, we set the size of Syntriever and all baseline models to (approximately) 125M. In alignment stage of Syntriever, we set $K=5$ by default, and set $N={K\choose 2}=10$. Detailed hyperparameters are in Appendix D.
@@
3.2 Main Results
We first present the supervised fine-tuning results on seven datasets for the retrieval task which are BeIR benchmarks with training datasets. The results are shown in Table 1. Compared to the second-best models, Syntriever improves the retrieval performances by: 18.6% on MSMARCO, 5.9% on HotpotQA, 2.5% on FiQA, 1.8% on SciFact, 8.3% on NFCorpus, and 4% on NQ. The base encoder for Syntriever is a pre-trained E5; still, Syntriever achieves performance gain over E5 by: 18.6% on MSMARCO, 10.7% on HotpotQA, 10.6% on FiQA, 9.2% on SciFact, 20.9% on NFCorpus, 5.6% on Fever and 5.7% on NQ. This shows that Syntriever can successfully distill LLMs’ capability into small retrievers and improve their performance by a large margin. Overall, Syntriever shows robust performances on datasets both in generalized and specialized domains. Our results show that small LMs can efficiently learn from the teacher model through synthetic data and can be successfully aligned through feedback, even without access to the output probability of black-box LLMs.
@@ 3.2 Main Results
3.3 Ablation study
We conduct an ablation study on Syntriever. We add or remove model components, and the effects on the performance are shown in Table 3. The results show that both synthesized query $(q^{\text{cot}})$ and passages $(p^{+},p^{-})$ in the distillation stage improve the retrieval performances. Overall, the distillation stage achieves an average gain of 8.2% over the base retriever. Results show that the retriever successfully learns from the parametric knowledge of LLMs during the distillation stage. Also, the alignment component $(c^{+}\succ c^{-})$ in Table 3 is shown to achieve performance gains of up to 8.8%. Our results show that the alignment component is significant for retrieval performance, considering that nDCG@K is sensitive to the fine-grained ranking of relevant passages.
@@ 3.3 Ablation study
3.4 Performances with different encoders
Syntriever is a framework for training encoders for retrieval, and thus can be combined with different sentence encoders. We experiment with various well-known encoders, e.g., ColBERT, SBERT, and Contriever, as the base encoders for Syntriever. Syntriever improves the performance by a large margin in all three retrieval models. The performance improvement is particularly high in nDCG@1 which concerns retrieving the exact passage relevant to the query. This is because the alignment stage in Syntriever helps the retriever with a fine-grained ranking of highly relevant passages. Overall, the results show that Syntriever is generally applicable to, and improves the performances of, various retrievers.
@@ 3.4 Performances with different encoders
3.5 Effects of Re-labeling Hallucination Passages
Table 5 shows that LLM self-verification and re-labeling are effective for the synthetic training by Syntriever. The performance improvement of self-verification on HotpotQA is relatively greater than other datasets. We found that, approximately 15% of synthetic positive passages were relabeled as hallucinations in the case of HotpotQA, whereas the proportion was about a few percent in other datasets. This indicates that the performance improvement through relabeling is likely higher for HotpotQA. In conclusion, removing hallucinations (and even re-using them as hard-negatives as in Syntriever) through self-verification is important for data synthesis, which is the case for most tasks utilizing LLMs Weng et al. (2023).
@@ 3.5 Effects of Re-labeling Hallucination Passages
3.6 Weaker but Cheaper LLMs can be effective
We examine how the LLM capabilities affect distillation and alignment performances. We consider two LLMs: GPT-4o vs. GPT-4o-mini, where GPT-4o is the larger and more capable model. First, we compare the distillation capabilities of two LLMs. Table 6 shows the comparison, where Syntriever is trained only up to the distillation stage. Interestingly, the distillation performance of GPT-4o-mini is better than GPT-4o on NFCorpus. Considering the datasets concern different knowledge domains (SciFact: scientific, NFCorpus: medical), smaller models may be better at teaching than larger ones in certain domains. Our results interestingly coincide with recent findings that weaker models may be better at teaching than stronger models in domains like math problem solving Bansal et al. (2024). Next, we examine the alignment capabilities of LLMs. For a fair comparison, two models are first trained by GPT-4o in the distillation stage, and then trained by different LLMs in the alignment stage. Table 7 shows that the larger model (GPT-4o) is better at alignment. It is challenging to rank top-$K$ passages retrieved by a distilled retriever, requiring a deep understanding of various contexts, and thus larger models may be more favored for the task. Overall, smaller models appear to be quite competitive, i.e., the performance gap is small or even better in some domains. Thus, our prospect is that distillation/alignment through small models will become an increasingly good alternative, especially under a fixed compute budget Bansal et al. (2024).
Method FiQA NFCorpus SciFact Partial Plackett Luce 41.9 43.3 80.5 Bradley Terry 35.7 36.1 79.8
Table 8: Comparison of preference modeling. Results are in nDCG@10.
3.7 Comparison of Preference Modeling Methods
3.8 Effects of the number of retrieved passages during alignment
3.9 Quality of Synthetic Positives
Passages Can derive answer Cannot derive answer Synthetic positive 88% 12% Ground truth 84% 16%
@@ Appendix A Reproducibility Statement
Black-box LLMs. We experiment with GPT-4o and GPT-4o-mini for synthetic data generation. Those models are accessible by OpenAI API. We generate LLMs’ responses using our prompt templates in Appendix C.
@@
Appendix B Derivation of (12)
@@
Appendix C Prompt templates
C.1 Prompt template of positive passage generation $(\mathcal{P}_{+})$.
@@ Appendix C Prompt templates
C.2 Prompts for generating plausible but irrelevant passages $(\mathcal{P}_{-})$.
@@
C.3 Prompt template of relabeling for synthetic passages $(\mathcal{P}_{\text{Relabel}})$.
@@
C.4 Prompts for pair-wise comparison of two passages $(\mathcal{P}_{\text{Compare}})$.
@@ Appendix D Hyperparameters
Appendix E Dataset Statistics